Molten aluminium chloride when electrolysed using 0.1F deposits x gram equivalent of aluminium. Find x in g equivalent : 1. 0.3 g equivalent 2. 0.9 g equivalent 3. 0.1 g equivalent 4. 0.6 g equivalent

Molten aluminium chloride when electrolysed using 0.1F deposits x gram equivalent of aluminium. Find x in g equivalent : 
1. 0.3 g equivalent
2. 0.9 g equivalent
3. 0.1 g equivalent
4. 0.6 g equivalent

Grade:11

1 Answers

silky srivastava
5 years ago
Here in this question the charge used to electrolyse aluminum is 0.1F. 
The formula is:
w=ZQ
Where w is the mass (in grams) of aluminum electrolysed and Q is the charge used that is equal to 0.1F that is equal to 0.1× 96500 C.
Z is a constant that is = molecular mass of the metal /(charge on the metal ×96500)
So in this case Z= molecular mass of aluminum /( charge on aluminum×96500) = 27/( 3×96500) = 9.33×10^5
w= Z×Q= (9.33×10^5 )×( 0.1× 96500) = 0.9g
But we have to give answer in gram equivalent so,
Gram equivalent of 0.9 g aluminium = 0.9/charge on aluminium = 0.9/3 = 0.3 gram equivalent.